OpenStudy (anonymous):

(In the form of 0*infinity) Find the limit as x approaches pi/2 from the negative of (pi/2-x)secx

OpenStudy (anonymous):

\[\lim_{x \rightarrow \frac{ \pi }{ 2 }-}\left( \frac{ \pi }{2 }-x \right)\sec x\] \[put~x=\frac{ \pi }{2 }-h,h>0,h \rightarrow0~asx \rightarrow \frac{ \pi }{ 2 }-\] \[\lim_{h \rightarrow 0}\left\{ \frac{ \pi }{2 }-\left( \frac{ \pi }{ 2 }-h \right) \right\}\sec \left( \frac{ \pi }{ 2 }-h \right)\] \[\lim_{h \rightarrow 0}h \csc h=\lim_{h \rightarrow0}\frac{ h }{ \sin h }=?\]

OpenStudy (anonymous):

Why does limh→0{π2−(π2−h)}sec(π2−h) turn into limh→0hcsch? Was this done with differentiation?

OpenStudy (anonymous):

\[\sec \left( \frac{ \pi }{ 2 }-h \right)=\frac{ 1 }{\cos \left( \frac{ \pi }{ 2 }-h \right) }=\frac{ 1 }{\cos \frac{ \pi }{ 2 } \cos h+\sin \frac{ \pi }{2 }\sin h} \] \[\cos \frac{ \pi }{2 }=0,0*\cos h=0,\sin \frac{ \pi }{ 2 }=1\]

OpenStudy (anonymous):

\[=\frac{ 1 }{ \sin h }=\csc h\]

OpenStudy (anonymous):

general rule to remember\[\frac{ \pi }{ 2 }-x~ is~add ~co~if~absent,~delete~co~if~present\] for example sine in cosine,here co is added. secant to cosecant,here again co is added. cotangent to tangent,here co is deleted. cosine to sine ,here co is deleted.

OpenStudy (kc_kennylau):

That means: sine(pi/2-x) = cosine(x) tangent(pi/2-x) = cotangent(x) secant(pi/2-x) = cosecant(x)

OpenStudy (kc_kennylau):

And vice versa

OpenStudy (anonymous):

correct